# Angular Momentum Drain: Despinning Embedded Planetesimals

### v1

@@ 2.7 The Angular Momentum Transfer Efficiency
When averaging over a distribution of impacts that lack angular momentum with respect to the planetesimal, the projectile momentum term in equation 34 can be neglected as it averages to zero. We only need to calculate the rightmost term that depends on the momentum of escaping ejecta as seen in the rotating frame. Because the planetesimal is orbiting the star, it is likely to feel a headwind coming from different directions in the orbital plane (see Figure 1). We assume that $\Delta{\bf L}_{a}$ components that are not aligned with the planetesimal’s spin axis average to zero. This implies that the average change in angular momentum per impact for a distribution of projectiles is
@@
$\epsilon_{L}\equiv-\frac{\langle\Delta{\bf L}_{a}\rangle}{L_{a}}\frac{M_{a}}{m_{pj}}$ (36)
@@
Table 1: Planetesimal fiducial parameters assumed throughout the simulations unless other values are stated. These parameters are based on the Kuiper Belt Object (486958) Arrokoth (Keane et al., 2022; Brisset et al., 2022), with the radius being the spherical equivalent from the mass and density of Arrokoth. The projectile velocity is taken to be a value near the headwind velocity experienced by planetesimals in the outer disk (Quillen et al., 2024). For these fiducial parameters, the escape velocity from the planetesimal is 3.6 m/s.
@@
$C_{1,N}$ $k_{N}$ $\nu$ $n_{1}$ $n_{2}$ $H_{2}$ $p$ $q$
0.55 0.3 0.4 1.2 1.0 0.4 0.3 0.2
Table 2: Crater scaling coefficients are taken from Table 3 by Housen and Holsapple (2011). We adopt parameters typical of granular materials.
@@ 2.8 Gravitational Focusing
For higher mass planetesimals, those with escape velocity above the projectile velocity, the gravitational field of the planetesimal causes gravitational focusing. In this case the velocity and angle of impact is sensitive to the orbit of the projectile when approaching the planetesimal. We compute the velocity and location of the impact using a hyperbolic orbit, neglecting gas drag. To collide with a spherical object of mass $M_{a}$ and radius $R_{a}$, a projectile must have impact parameter $b$ lower than
@@ 2.9.2 Planetesimal Shape and Structure
For simplicity, we have neglected the shape of the planetesimal by assuming that it is spherical. If we allow for non-spherical planetesimals, the gravity vector at a surface location is not necessarily perpendicular to the surface, so the surface can have a slope with respect to the direction of self-gravity. Studies measuring crater volume, diameter, and axis ratio of oblique impacts into inclined granular media find that ejecta distributions also depend on surface slope (Takizawa and Katsuragi, 2020; Omura et al., 2021). In future, models could be extended to take into account impact angle, surface slope and the direction and values of ejecta velocity vectors that allow ejecta to escape. Such models could find an analogies to the Yarkosky effect, causing orbital drift, but mediated by momentum exchange with a distribution of projectiles.
@@ 3 Ejecta Distributions for Single Impacts
Figure 5: We show a polar plot of an impact with projectile density $\rho_{pj}=1000$ kg m-3 occurring at a Latitude = 0, Longitude = 0. We define Longitude = 0 as the furthest point on the headwind-facing hemisphere. (a) In the frame moving with the surface, we show ejecta velocity in color as a function of ejecta launch position and azimuthal angle. The radius on the plot refers to distance from impact point (in units of projectile radius) and the polar angle on the plot is $\zeta$, the azimuthal angle. Due to the rotation of the planetesimal, the impact angle in the frame rotating with the surface of this collision is $\theta_{I}\approx 87.1^{\circ}$ despite being a head-on collision. (b) Similar to Figure 5a, except we only show pixels that are above the escape velocity, highlighting regions of the crater that lead to angular momentum transfer and mass loss. Despite being a head-on collision, planetesimal rotation causes an asymmetric ejecta curtain.
@@
For each impact, we generate a 2 dimensional polar grid of ejecta launch positions ($x$ values) and azimuthal angles ($\zeta$ values). Here $x$ is the radial position within the crater centered at the site of impact impact (see Figure 3), in units of the projectile radius, $a_{pj}$. At each point on this grid we compute an ejecta velocity velocity in the rotating frame, an ejecta velocity vector in the inertial frame, and the ejecta mass. We calculate the crater radius $R_{cr,N}$ for a normal impact using Equation 14. Following the point-source approximation, we reject locations with $x<n_{1}a_{pj}$ or $x\leq n_{2,N}R_{cr}$ by setting the ejecta mass and velocity at those grid points to zero. For all remaining $x$ and $\zeta$ points within our grid, we calculate the ejecta velocity and mass. The ejecta velocity $v_{ej}$ in the rotating frame within an oblique impact crater is given by Equation 24, and crater radius $R_{cr}$ in equation 22.
@@ 3.2 Caveat: Ejecta Models
Cohesion is incorporated into a strength parameter describing the asteroid surface. However, this description may not capture granular dynamics, particularly if projectiles are similar in size to particles on the planetesimal surface.
@@ 4 Torque from a Distribution of Impacts
In the previous section we integrated the ejecta mass and velocity distributions for a single impact. In this section we integrate a distribution of impacts to estimate the total mass ejected and the angular momentum transferred to the planetesimal by a population of impactors.
@@
Figure 7: Mass loss ratio and angular momentum transfer efficiency for a planetesimal with the fiducial values from Table 1 as a function of planetesimal radius on the x-axis. In each panel pair (a-b, c-d, e-f, g-h), the effect of varying a single parameter is shown. Panels (a,b) varies $\pi_{4}$, (b,c) varies planetesimal strength $Y_{a}$, (e,f) varies the distant headwind velocity $u_{hw}$, and (g,h) varies the planetesimal spin rate. The left column (subfigures a,c,e,g) show the angular momentum transfer efficiency and the right column (subfigures b, d, f, h) show the mass loss ratio. In the mass loss ratio plots b, d, f, h, we indicate regions of net accretion in green shading and net erosion in red shading. We find that the greatest influence on the angular momentum transfer are due to $\pi_{4}$ (the ratio of planetesimal to projectile density) and headwind velocity $u_{hw}$, while the mass loss ratio is most affected by the planetesimal strength $Y_{a}$ and head wind velocity $u_{hw}$.
@@
In Figure 7, we plot the angular momentum transfer efficiency, $\epsilon_{L}$ and the mass loss ratio $\epsilon_{M}$ as a function of planetesimal radius $R_{a}$. All quantities, when not specified, are taken to be those of the fiducial parameters listed in Table 1. In Figure 7a and 7b, we vary the projectile density $\rho_{pj}$ giving different $\pi_{4}$ values. These panels show how the density ratio between the projectile and planetesimal affect the mass loss ratio and angular momentum change as a function of planetesimal radius $R_{a}$. Figures 7c - 7h show how variations in strength $Y_{a}$, headwind velocity $u_{hw}$, and spin $\Omega_{a}$ affect the mass loss ratio and angular momentum transfer efficiency.
@@
The panels in the left column of Figure 7 show that the angular momentum transfer efficiency is positive, confirming that the mechanism proposed by Dobrovolskis and Burns (1984) operates in a low velocity regime. However, the angular transfer efficiency is low, at most $\epsilon_{L}\sim$ 0.07 in the most extreme case at $\pi_{4}=0.25$. We find that the mechanism for spin down is inefficient. The panels in the right column show that, for our fiducial parameters, the total ejected mass remains in an accretionary regime, with the accretion being more efficient for larger radii planetesimals which have higher escape velocities. Thus the angular momentum drain mechanism proposed by Dobrovolskis and Burns (1984) also operates in an accretionary impact regime.
@@
In Figure 7, for planetesimal radius $R_{a}\lesssim 10$ km, the angular momentum transfer efficiency $\epsilon_{L}$ is nearly independent of planetesimal radius. Above $R_{a}\sim 10$ km the angular momentum transfer efficiency tends to decrease, and this we attribute to due to gravitational focusing increasing the projectile impact velocity.
In Figure 7c - 7d we see low-strength planetesimal materials exhibit marginally increased angular momentum transfer efficiency at large planetesimal radii. At lower mass and radii, the effect flips, with low strength planetesimals exhibiting the least angular momentum spindown, with an intermediate strength values being most effective in this regime. At all points, especially at lower mass and radii, lower strength bodies exhibit increased mass loss because each crater sheds more material from the planetesimal when compared to one on a higher strength planetesimal.
@@
In Figure 7e-7f we show how the headwind velocity affects the mass loss ratio and angular momentum transfer efficiency. High wind velocities result in an increase in mass loss ratio, $\epsilon_{M}$, and a decrease in angular momentum transfer efficiency, $\epsilon_{L}$. At low headwind velocities, the ejecta velocity is smaller, thus planetesimal rotation has a greater influence in determining which ejecta escapes, as illustrated by Figure 4. This leads to an increase in escaping ejecta on the trailing hemisphere, leading to larger spin-down. At higher headwind velocities, the ejecta velocity is greater and we find more ejecta escaping downrange on the leading hemisphere, leading to a lower overall spin-down.
@@
In Figure 7g-7h, we show sensitivity to the planetesimal spin. We find that the accretion rate is insensitive to the spin rate, and the angular momentum transfer efficiency decreases as a function of decreasing spin rate.
For all but one variable we see the changes to the mass loss ratio and angular momentum transfer efficiency are aligned, we see an increase/decrease in both effects across all planetesimal masses. However, the headwind velocity exhibits the opposite trend, with an increase in $\epsilon_{L}$ associated with a decrease in $\epsilon_{M}$. To explore why this is true, we create a new figure to investigate the asymmetric influence of planetesimal rotation on the angular momentum transfer efficiency for the leading vs trailing hemispheres.
@@
Among the varied parameters $\pi_{4},Y_{a},u_{hw},\Omega_{a}$ Figure 7a shows that the angular momentum transfer is most sensitive to the ratio of planetesimal to projectile density, $\pi_{4}$. A high density projectile colliding with a low density planetesimal will yield significantly more ejecta and so more angular momentum transfer than when the densities are equal.

### v2

@@ 2.7 The Angular Momentum Transfer Efficiency
When averaging over a distribution of impacts that lack angular momentum with respect to the planetesimal, the projectile momentum term in equation 34 can be neglected as it averages to zero. We only need to calculate the rightmost term that depends on the momentum of escaping ejecta as seen in the rotating frame. Because the planetesimal is orbiting the star, it is likely to feel a headwind coming from different directions in the orbital plane (see Figure 1). We assume that $\Delta{\bf L}_{a}$ components that are not aligned with the planetesimal’s spin axis average to zero. Within a crater, the crater ejecta is symmetric about the uprange-downrange line, and thus the binormal components average to zero. The normal component of the crater ejecta do not average to zero due to asymmetry in the projectile impact velocity in the rotating frame. This would lead to variations in the radial position of the planetesimal orbit and is a consideration for future work. The spin angular momentum of the planetesimal would not be affected by this, but rather the orbital angular momentum. This implies that the average change in spin angular momentum per impact for a distribution of projectiles is
@@
$\epsilon_{L}\equiv-\frac{|\langle\Delta{\bf L}_{a}\rangle|}{L_{a}}\frac{M_{a}}{m_{pj}}$ (36)
@@
Table 1: Planetesimal fiducial parameters assumed throughout the simulations unless other values are stated. These parameters are based on the Kuiper Belt Object (486958) Arrokoth (McKinnon et al., 2020; Keane et al., 2022; Brisset et al., 2022), with the radius being the spherical equivalent from the mass and density of Arrokoth. McKinnon et al. 2020 give a minimum cohesive strength range of 100-400 Pa for Arrokoth determined as a minimum required strength for stabilization of the neck that Arrokoth exhibits. The projectile velocity is taken to be a value near the headwind velocity experienced by planetesimals in the outer disk (Quillen et al., 2024). For these fiducial parameters, the escape velocity from the planetesimal is 3.6 m/s.
@@
Regime Material $C_{1,N}$ $k_{N}$ $\nu$ $\mu_{N}$ $n_{1}$ $n_{2}$ $H_{1}$ $H_{2}$ $p$ $q$
Strength SFA 0.55 0.3 0.4 0.4 1.2 1.0 - 0.4 0.3 0.2
Gravity Sand 0.55 0.3 0.4 0.41 1.2 1.3 0.59 - 0.3 0.2
Table 2: Crater scaling coefficients are taken from Table 3 by Housen and Holsapple (2011) for SFA (sand/fly ash) and for sand. We adopt parameters typical of granular materials.
@@ 2.8 Gravitational Focusing
For higher mass planetesimals, those with escape velocity above the projectile velocity, the gravitational field of the planetesimal causes gravitational focusing. In this case the velocity and angle of impact is sensitive to the orbit of the projectile when approaching the planetesimal.
We compute the velocity and location of the impact using a hyperbolic orbit, neglecting gas drag. To collide with a spherical object of mass $M_{a}$ and radius $R_{a}$, a projectile must have impact parameter $b$ lower than
@@ 2.9.2 Planetesimal Shape and Structure
For simplicity, we have neglected the shape of the planetesimal by assuming that it is spherical. If we allow for non-spherical planetesimals, the gravity vector at a surface location is not necessarily perpendicular to the surface, so the surface can have a slope with respect to the direction of self-gravity. Studies measuring crater volume, diameter, and axis ratio of oblique impacts into inclined granular media find that ejecta distributions also depend on surface slope (Takizawa and Katsuragi, 2020; Omura et al., 2021). In future, models could be extended to take into account impact angle, surface slope and the direction and values of ejecta velocity vectors that allow ejecta to escape. Such models could find an analogies to the Yarkovsky effect, causing orbital drift, but mediated by momentum exchange with a distribution of projectiles.
@@ 2.9.4 Rotating Projectiles
3 Impact Regime
3.1 Gravity vs Strength Impact regime
Our numerical procedure for calculating the angular momentum transfer efficiency $\epsilon_{L}$ and fraction of mass per unit projectile mass $\epsilon_{M}$ depend on whether impacts are in the gravitational or strength regime through the crater radius on the regime (equations 13 and 14). The transition between gravity regime and strength regimes depends on the dimensionless $\pi_{2},\pi_{3},\pi_{4}$ parameters (Holsapple, 1993). The gravity regime is where
$\pi_{2}\gtrsim\left(\frac{H_{1}}{H_{2}}\right)^{(2+\mu_{N})/\mu_{N}}\pi_{4}^{\nu}\pi_{3}^{(2+\mu_{N})/2}$ (43)
(following Holsapple 1993; Housen and Holsapple 2011). We solve for the projectile velocity at the transition between regimes,
$v_{pj}\gtrsim\left(\frac{Y_{a}}{\rho_{a}}\right)^{\frac{2+\mu}{2\mu}}\left(\frac{\rho_{a}}{\rho_{pj}}\right)^{\frac{\nu}{\mu}}(g_{a}a_{pj})^{-\frac{1}{\mu}}.$ (44)
Because of the composition and structure of primordial materials are unknown, it is difficult to definitively determine the impact regime. The study of the ejecta curtain from the artificial impact on asteroid 162173 Ryugu (Arakawa et al., 2020), and the presence of ejecta near crater rims on asteroids Ryugu and 101955 Bennu (Arakawa et al., 2020; Perry et al., 2022) infer that the strength of the materials near the surface is below 1 Pa. The TAGSAM contact on Bennu gave a similar strength limit (Walsh et al., 2022). These strength limits lie below the $\sim 25$ Pa asteroid cohesion strength estimated from van der Waals contact forces (Scheeres et al., 2010; Sánchez and Scheeres, 2014).
[…]
In Figure 5 we plot lines delineating the transition between gravity and strength regimes condition with $\rho_{a}=235{\rm\ kg\ m^{-3}}$ and $\pi_{4}=1$ for different values of planetesimal and projectile radii. The gravity regime is at a higher velocity at lower surface gravity (smaller planetesimal radius), and smaller projectile radius. We find that for low impact velocity (between 1 and 100 m/s), low density, planetesimal radii from 1 to 100 km, and projectiles below 1 m in radius, the impacts would mostly be in the strength regime as long as $Y_{a}\gtrsim 10$ Pa. For comparison, Sánchez and Scheeres (2014) infer a cohesive strength of $Y_{a}\approx 25$ Pa for rocky rubble piles regardless of size from Van der Waals forces. In summary, we expect the collisions considered in this study to primarily occur within the strength regime, however due to the proximity to the transition between strength and gravity regime, the crater sizes using gravity regime or strength regime do not differ greatly.
3.2 Projectile Size
The size distribution of projectiles is relevant for determining the impact regime, as discussed in the previous section, but also determines the size of the change in angular momentum by each impact.
Particles concentrated by the Streaming instability are expected to be within $0.04\leq{\rm St}\leq 3$ (Rucska and Wadsley, 2023; Carrera et al., 2015). If particles experience compacted growth, are primarily porous and icy, or porous icy and experience erosion, growth rapidly occurs up until St =1, at which point inward drift begins to remove the material (Krijt et al., 2016). Thus, we may expect an upper limit for particles with St = 1 to be present within the streaming instability in the outer disk. Many impacts on young planetesimals are expected to be from objects that are within this range (Quillen et al., 2024), though late states of coalescence may involve impacts from larger objects (e.g., Nesvorný et al. 2019).
Particles with $St\sim 1$ have a radius of about 1 m (based on stopping time estimates for the disk model by Quillen et al. 2024). Based on stopping time arguments, we expect a minimum particle radius to be $\sim 10^{-3}$ mm (Quillen et al., 2024).
For projectile radii ranging from 1 cm to 1 m, assuming the fiducial values listed in Table 1 over a range of planetesimal radii, impacts occur in a region close to the transition between strength and gravity regimes. While in the strength regime, the angular momentum transfer and mass loss is not dependent on the projectile radius. Thus, 10 collisions with a projectile of mass 100 kg, and 1 collision with a projectile of mass 1000 kg, all other parameters being equivalent, would have the same net result on ejected mass and angular momentum transfer efficiency.
In the gravity regime, when projectile radius sets the $\pi_{2}$ parameter, we assume a projectile radius of $a_{pj}=1$ m, as the largest projectile radii and subsequently projectile mass will dominate the angular momentum transfer and mass loss. Decreasing the projectile radius results in lower planetesimal strength being required for impacts to occur in the gravity regime. As such, we choose an expected particle size with the largest Stokes number to ensure we remain in the same impact regime across all planetesimal radii tested.
@@ 4 Ejecta Distributions for Single Impacts
Figure 6: We show a polar plot of an impact with projectile density $\rho_{pj}=1000$ kg m-3 occurring at a Latitude = 0, Longitude = 0. We define Longitude = 0 as the furthest point on the headwind-facing hemisphere. (a) In the frame moving with the surface, we show ejecta velocity in color as a function of ejecta launch position and azimuthal angle. The radius on the plot refers to distance from impact point (in units of projectile radius) and the polar angle on the plot is $\zeta$, the azimuthal angle. Due to the rotation of the planetesimal, the impact angle in the frame rotating with the surface of this collision is $\theta_{I}\approx 85.3^{\circ}$ despite being a head-on collision. (b) Similar to Figure 6a, except we only show pixels that are above the escape velocity, highlighting regions of the crater that lead to angular momentum transfer and mass loss. Despite being a head-on collision, planetesimal rotation causes an asymmetric ejecta curtain.
@@
For each impact, we generate a 2 dimensional polar grid of ejecta launch positions ($x$ values) and azimuthal angles ($\zeta$ values). Here $x$ is the radial position within the crater centered at the site of impact (see Figure 3), in units of the projectile radius, $a_{pj}$. At each point on this grid we compute an ejecta velocity in the rotating frame, an ejecta velocity vector in the inertial frame, and the ejecta mass. We calculate the crater radius $R_{cr,N}$ for a normal impact using Equation 13 or 14, depending upon whether the impact is in the strength or gravity regime. Following the point-source approximation, we reject locations with $x<n_{1}a_{pj}$ or $x\leq n_{2,N}R_{cr}$ by setting the ejecta mass and velocity at those grid points to zero. For all remaining $x$ and $\zeta$ points within our grid, we calculate the ejecta velocity and mass. The ejecta velocity $v_{ej}$ in the rotating frame within an oblique impact crater is given by Equation 24, and crater radius $R_{cr}$ in equation 22.
@@ 4.2 Caveat: Ejecta Models
Cohesion is incorporated into a strength parameter describing the asteroid surface. However, this description may not capture granular dynamics, particularly if projectiles are similar in size to particles on the planetesimal surface. Cohesion is often described via an energy per unit area, so it is likely to more strongly affect smaller particles as they have a large surface area relative to their mass. Experimental studies find that micron sized water-ice particles stick below a threshold velocity of 9.6 m s-1 (Gundlach and Blum, 2015).
@@ 5 Torque from a Distribution of Impacts
In the previous section we integrated the ejecta mass and velocity distributions for a single impact. In this section we integrate a distribution of 10,000 impacts to estimate the total mass ejected and the angular momentum transferred to the planetesimal by a population of low mass impactors.
@@
Figure 8: Mass loss ratio and angular momentum transfer efficiency for a planetesimal with the fiducial values from Table 1 as a function of planetesimal radius on the x-axis. In each panel pair (a-b, c-d, e-f, g-h), the effect of varying a single parameter is shown. Panels (a,b) varies $\pi_{4}$, (b,c) varies planetesimal strength $Y_{a}$, (e,f) varies the distant headwind velocity $u_{hw}$, and (g,h) varies the planetesimal spin rate. The left column (subfigures a,c,e,g) show the angular momentum transfer efficiency and the right column (subfigures b, d, f, h) show the mass loss ratio. In the mass loss ratio plots b, d, f, h, we indicate regions of net accretion in green shading and net erosion in red shading. We find that the greatest influence on the angular momentum transfer are due to $\pi_{4}$ (the ratio of planetesimal to projectile density) and headwind velocity $u_{hw}$, while the mass loss ratio is most affected by the planetesimal strength $Y_{a}$.
@@
In Figure 8, we plot the angular momentum transfer efficiency, $\epsilon_{L}$ and the mass loss ratio $\epsilon_{M}$ as a function of planetesimal radius $R_{a}$. All quantities, when not specified, are taken to be those of the fiducial parameters listed in Table 1. In Figure 8a and 8b, we vary the projectile density $\rho_{pj}$ giving different $\pi_{4}$ values. These panels show how the density ratio between the projectile and planetesimal affect the mass loss ratio and angular momentum change as a function of planetesimal radius $R_{a}$. Figures 8c - 8h show how variations in strength $Y_{a}$, headwind velocity $u_{hw}$, and spin $\Omega_{a}$ affect the mass loss ratio and angular momentum transfer efficiency.
@@
The panels in the left column of Figure 8 show that the angular momentum transfer efficiency is positive, confirming that the mechanism proposed by Dobrovolskis and Burns (1984) operates in a low velocity regime. The panels in the right column show that, for our fiducial parameters, the total ejected mass remains in an accretionary regime for larger radii planetesimals which have higher escape velocities. At small planetesimal radii the total ejected mass is in a transitional region between erosional and accretionary dependent on the specifics of the projectile/material properties. Thus the angular momentum drain mechanism proposed by Dobrovolskis and Burns (1984) also operates in an accretionary impact regime.
@@
In Figure 8, for planetesimal radius $R_{a}\lesssim 5$ km, the angular momentum transfer efficiency $\epsilon_{L}$ is nearly independent of planetesimal radius, with Figure 8c showing a clear deviation from this overall trend, which we better describe below and in Section 5.4. Above $R_{a}\sim 5$ km the angular momentum transfer efficiency tends to remain flat or slightly decrease, with the exception of low headwind velocity increasing angular momentum transfer due to gravitational focusing accelerating the projectile prior to impact.
In Figure 8a - 8b we show for a range of density ratios, with the planetesimal or projectile being 4x the density of the other, with equal densities in the middle. At small planetesimal radii, we see a trend of spinup occurring for large values of $\pi_{4}$. This effect is discussed in greater detail when discussing the variation in planetesimal strength and in Section 5.4. We find that the mass loss ratio is greatest when the projectile density is less than the planetesimal for low-radii planetesimals. This is due to the mass loss ratio being the mass ejected in units of projectile mass. Decreasing the projectile mass while leaving the mass ejected unchanged will lead to a greater mass loss ratio. This is specific for low-radii planetesimals where we find a majority/all of the crater mass escapes which would occur for low-strength, small-radii planetesimals. As such, the change in projectile mass is a primary driver in the increased mass loss ratio in these cases. At larger planetesimal radii, we find that a more dense projectile results in greater a greater mass loss ratio.
In Figure 8c - 8d we find that strengthless planetesimals exhibit a much greater angular momentum transfer than planetesimals with low strength. If planetesimal strength is important in crater scaling, small variations in the planetesimal strength do not lead to large changes in the angular momentum transfer. Angular momentum drain is most effective when the cratering occurs in the gravity regime. We see a dip in the angular momentum transfer and a reversal to spin-up in the strengthless material case. This dip can begin to be seen in the low-strength case as well, showing as the small dip at $\sim 5$ km. Additionally, this dip is the same as is seen for $\pi_{4}>1$ in Figure 8a, $u_{hw}>25$ m/s in Figure 8e, and all $\Omega_{a}$ in Figure 8g. This effect occurs due to highly oblique impacts in low-strength material exhibiting a butterfly crater effect, seen for oblique impacts by Gault and Wedekind 1978 and Luo et al. 2022, for the high-velocity ejecta. Impacts on the spin-down hemisphere see more of a butterfly effect in the escaping ejecta compared to an equivalent location impact on the spin-up hemisphere. This results in reduced angular momentum transfer in the spin-down direction, causing a decrease in net spin-down, or a complete reversal to spin-up at the lowest strength. This effect may or may not be physical, and is further discussed in Section 5.4. At all points, especially at lower mass and radii, lower strength bodies exhibit increased mass loss because each crater sheds more material from the planetesimal when compared to one on a higher strength planetesimal.
@@
In Figure 8e-8f we show how the headwind velocity affects the mass loss ratio and angular momentum transfer efficiency. Increases in the headwind velocity result in increased mass loss and angular momentum transfer. As the headwind velocity decreases, for small radii planetesimals the angular momentum transfer approaches 0, however as the planetesimal radius increases, gravitational focusing accelerates the projectiles, increasing the impact velocity. Even at low headwind velocities and planetesimal radii where gravitational focusing is not relevant, the angular momentum transfer can seem quite large. At these low headwind velocities, the ejecta velocity is smaller, thus planetesimal rotation has a greater influence in determining which ejecta escapes, as illustrated by Figure 4. This leads to an increase in escaping ejecta on the trailing hemisphere, leading to larger spin-down. While this effect ensures that low headwind velocities can result in angular momentum drain, the decrease in ejecta velocity and leads to no ejecta escaping and thus a total shutdown of angular momentum transfer due to crater ejecta.
@@
In Figure 8g-8h, we show sensitivity to the planetesimal spin. We find that the accretion rate is insensitive to the spin rate, and the angular momentum transfer efficiency decreases as a function of decreasing spin rate.
@@
Among the varied parameters $\pi_{4},Y_{a},u_{hw},\Omega_{a}$ Figure 8a shows that the angular momentum transfer is most sensitive to the ratio of planetesimal to projectile density, $\pi_{4}$. A high density projectile colliding with a low density planetesimal will yield significantly more ejecta at high velocity and thus more angular momentum transfers than when the densities are equal.